14. How much would you have to deposit each quarter at 6% compounded quarterly, in order to have $32,000 in 5 years?
Answer #1

Answer: for calculating how much to deposit each quarter, following formule is used: nxt А A = P PE 1+ a) n n where, А A = Am32000 x 0.015 =P (11346855 - 1) 480 i) p = 0.34685 =) p = $ 1383.86 Sol $1383. 86 should be deposited each quartes in order t

Thankyou.......